A partnership has the following capital balances:

  Arlo (40% of gains and losses) $ 62,000
  Band (25%) 70,000
  Carlyle (35%) 144,000

David is going to invest $54,000 into the business to acquire a 30 percent ownership interest. Goodwill is to be recorded. What will be David’s beginning capital balance?

a. $54,000.

b. $99,000.

c. $82,800.

d. $118,286.
Answer #1
Total Capital before investment 276000 =62000+70000+144000
David's investment = 30% (Original capital plus David's investment)
54000+Goodwill = 0.30*(276000+54000+Goodwill)
54000+Goodwill = 99000+0.30 Goodwill
0.70 Goodwill = 99000-54000
0.70 Goodwill = 45000
Goodwill = 45000/0.70 = 64286
David’s beginning capital balance = 54000+64286= $118,286
Option D $118,286 is correct